OpenStudy (anonymous):

# of jobs in 2004 is 80 # of jobs in 2014 is expected to be 172 what is the percent increase show work please:)

OpenStudy (ash2326):

Percentage Increase in Jobs \[=\frac{\text{# of jobs in 2014}-\text{# of jobs in 2004}}{\text{# of jobs in 2004}} \times 100\] So we have \[\frac{172-80}{80}\times 100=???\] Can you find now?

OpenStudy (lgbasallote):

\[\large \frac{172-80}{80} \times 100\] i believe that is the formula

OpenStudy (anonymous):

53.49 but the answer is 115%

OpenStudy (lgbasallote):

did you do 172 - 80 right?

OpenStudy (lgbasallote):

should be around 92

OpenStudy (ash2326):

Check again, you'll get the answer?

OpenStudy (anonymous):

nvm lol:)